### v1

@@ 6 CONCLUSIONS
To validate and compare these advanced architectures, we developed a unified link analysis framework that provides a high-fidelity assessment by incorporating unique environmental factors. Second, building on this framework, we assessed the reliability via outage probability to quantify the operational robustness of architectures under dynamic conditions. Through these, our work shows the advantages of NSN integrated architectures but also underscores the emerging requirement to build the resilient, dynamic, and high-performance communication backbone. Therefore, we proposed the inter-domain space digital twin, a dynamic decision-making engine that autonomously analyzes the networks in real-time to dynamically select the most reliable and power-efficient communication path. Final results show the proposed approach ensures optimal performance and resource utilization for the next generation of sustained lunar exploration.

### v2

@@ 6 CONCLUSIONS
To validate and compare these advanced architectures, we developed a unified link analysis framework that provides a high-fidelity assessment by incorporating unique environmental factors. Second, building on this framework, we assessed their reliability via outage probability to quantify the operational robustness of architectures under dynamic conditions. Through these, our work shows the advantages of NSN integrated architectures but also underscores the emerging requirement to build a resilient, dynamic, and high-performance communication backbone. Therefore, we proposed the inter-domain space digital twin, a dynamic decision-making engine that autonomously analyzes the networks in real-time to dynamically select the most reliable and power-efficient communication path. Final results show that the proposed approach ensures optimal performance and resource utilization for the next generation of sustained lunar exploration.
{IEEEbiography}[
]Selen Geçgel Çetin (Member, IEEE) received the B.S. degree from Yildiz Technical University in 2016, and the M.S. and Ph.D. degrees from Istanbul Technical University (ITU) in 2019 and 2025, respectively. She is a member of the Wireless Communication Research Laboratory at ITU and the Poly-Grames Research Center at Polytechnique Montréal. Her research focuses on the intersection of machine learning and wireless communications, with a specific emphasis on intelligent designing and adaptive managing for secure and reliable space communication.
{IEEEbiography}[
]Barış Dönmez (Graduate Student Member, IEEE) received the B.Sc. and M.Sc. degrees (with high honors) in electrical and electronics engineering, in 2009 and 2022, respectively, from FMV Işık University, Istanbul, Turkiye. He is currently pursuing his Ph.D. degree in electrical engineering at Polytechnique Montréal, Montréal, QC, Canada. His research interests include communication and energy harvesting systems in space networks.
{IEEEbiography}
] Güneş Karabulut Kurt [M’00, SM’14] is a Canada Research Chair (Tier 1) in New Frontiers in Space Communications and Professor at Polytechnique Montréal, Montréal, QC, Canada. She is also an adjunct research professor at Carleton University, ON, Canada. Güneş is a Marie Curie Fellow and has received the Turkish Academy of Sciences Outstanding Young Scientist (TÜBA-GEBIP) Award in 2019. She received her Ph.D. degree in Electrical Engineering from the University of Ottawa, ON, Canada.